Fermi National Laboratory


Accelerator Update

Wednesday December 3
The day shift began with Tevatron (TeV) experts conducting studies, with MiniBooNE taking beam, and the Antiproton Source (Pbar) stacking.

A Linac expert conducted ten minutes of H- Source studies at 10:22 AM.

Around noon, Recycler took shots of antiprotons from Pbar. Unfortunately, 30mA of antiprotons went into the first shot, which wasn't intended. And worse, the shot didn't make it into the Recycler due to a RF mismatch. Experts made adjustments and continued. The last shot was sent at 1:47 PM.

At 4 PM, the TeV suffered a sector D1 quench due to the low beta magnets tripping off. Pbar resumed stacking.

Operators began preparing the Collider Detector Facility (CDF) and D-zero (D0) experimental halls ready for access at 4:23 PM.

D0 was out of access by 6 PM. CDF never went in on access.

Operations put the TeV into shot setup at 6:32 PM.

At 6:43 PM, a Booster kicker tripped off. An expert came in and had the faulty cable repaired by 8 PM.

At 9 PM, some Pbar and Main Injector (MI) problems were delaying the shot setup. Operations and experts worked at clearing these problems.

A Linac RF station (LRF5) tripped off at 11:34 PM due to an airflow bit problem. Operators called in a Linac technician.

Thursday December 4
The midnight shift began with the TeV shot setup on hold for LRF5 repairs.

A Linac tech had LRF5 back in service by 2:30 AM.

Operations restarted the TeV shot setup at 2:38 AM and at 4:24 AM established store 3080. Its initial luminosity was 2.2E31. Pbar resumed stacking.

At 6:28 AM, a skew quadrupole correction circuit (T:SQ) tripped off and caused store 3080 to abort. There was no quench. This device was back in service by 11 AM.

Operators made the D0 and CDF collision halls ready for access at a little after 7 AM. They were both out of access by 9:15 AM.

Operations put the TeV into shot setup at 11 AM.

At 12:41 PM, the TeV quenched at sector A4.

Operations put the TeV through a dry squeeze and then into shot setup at 3:07 PM.

At 5:40 PM, Operations established store 3083 with an initial luminosity of 1.9E31. Pbar resumed stacking.

A Pbar stacking problem developed at 9:45 PM. Operators consulted with an expert and then switched to a different RF cavity (ARF1-2) for stacking. This solved the problem.

Friday December 5
The midnight shift began with Pbar stacking, with Operations monitoring store 3083, and with MiniBooNE taking beam.

At 4 AM, Operations turned off beam for MiniBooNE to investigate a possible problem with their Horn. Operators called an expert, but he couldn't find any problems. Operations restored beam about thirty minutes later.

Today's Plans
Operations will terminate the present store and put the TeV into shot setup around noon today. Recycler experts will take shots of antiprotons that add up to a total of 20mA later this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/5/2003   email Fermilab